# DrawTalking: Building Interactive Worlds by Sketching and Speaking

### v3

@@ 4.6.2. Processing Steps
The system receives language input and outputs a directed graph data structure encoding the dependencies between and semantic roles of words.
Our implementation uses the Spacy (v 3.1.4)(Montani et al., 2023) library to output a dependency tree per-sentence, along with a library called ”coreferee” to fill-in coreference information.
@@
Lastly, the application traverses the structure and generates a final application-defined structure S3 that it can evaluate – in our case, a mix between scriptable virtual machine and animation engine akin to (Perlin and Goldberg, 1996; Resnick et al., 2009) or a runnable node-based program (cycling74, 2018). S3 can call these scripts and retain S2 for reference. When traversing nodes in S2 labeled ACTION (the verbs) we lookup the appropriate previously-defined script for that action and insert a reference to it in S3. The arguments for that action, i.e. semantic role keys mapped to object ids or types, are inserted into a lookup table specific to the ACTION script instance so when the action executes, it knows what objects to modify. Loops and timed waits are also inserted into S3. The application evaluates the final structure according to its own interpretation. This results in any potential application-specific side-effect – in our case, animation, simulation, rule creation, application state-changes and so on.
@@
Figure 12. Concrete Client/Server Details. Text data are sent to a server for natural language processing over a data channel and forwarded to an NLP-focused sub-process on the same machine. The semantic parse data return to the client. A command channel handles user events independently (e.g. discarding commands interrupts in-flight stale NLP).
@@
[A system diagram for the client-server architecture and data flow]A system diagram for the client-server architecture and data flow. The user is shown as a stick-figure with a quote bubble reading ”speech” and the text ”pen+touch” both connected to the main DrawTalking client node in the system diagram (indicating possible sources of user input.) The client node contains a real-life photo of the researcher’s iPad Pro used for the prototype, atop a blue square with the name ”DrawTalking” on it, as well as a small cartoon microphone icon. Below the square reads: ”Main C++, C; Platform: Obj C, Swift, Metal Graphics; On-Device Speech Recognition”. 3 arrows connect the client node to a server node. The arrows are referred to as ”Wireless Local Connection” (to the server. Arrow 1 (outgoing) is for text data. Arrow 2(incoming) is for the semantics parse. Arrow 3 is for control messages to the server; In the server node is a real-life photo of the Macbook Pro used for the server, above two nodes: left: the control server labeled NodeJS; right the Language Processing Sub-Process labeled Python 3. 2 arrows between server nodes (both surrounded by an enclosing rectangle): arrow 1: (outgoing) is for text. arrow 2: (incoming) is for the resulting semantics parse, which is sent back to the control server and back to the DrawTalking client along the other arrow incoming to the DrawTalking client.
@@ 6. Results and Discussion
All participants understood and learned the mechanics, and produced thoughtful discussion. They tried their own ideas for how to make things or test their own understanding out of curiosity. All became comfortable with the core mechanics and some were excited to continue playing with the tool. They identified use cases include educational applications, rapid videogame and design prototyping, paper prototyping, user interface design, presentation, language learning, and visual-oriented programming.
@@ 6.2. Semantics-visuals correspondence is valued
For the semantics diagram, participants understood it as a helpful means to introspect and correct the system’s understanding of user input. P2: ”[The diagram] is pretty important. Especially if you’re getting to used to the program, you might say something wrong and it interprets it in a different way and you can correct it.” P8: ”This is a really cool debug thing. I immediately understand what happens when I read it.” Additionally, the interface simplifies the input, making it easier to read: P7: ”[It’s] showing what it’s going to do. It’s taking the verbs and cutting all the other stuff out. The feedback suggests that the diagram is valuable for instilling confidence in the system, and that the interface achieved D3.
For the text view, we observed users (e.g. P1, P5) habitually clearing the transcript using ”discard” on their own. We asked why and some participants reported wanting to keep the interface clean. We did not foresee this, but this user behavior serves a dual-purpose: the user is aware of what they say, which means they are perhaps likelier to catch errors, and they also help by keeping the system input from growing too long. We can take this as a positive: the system helps the user, and the user helps the system. We can also take this as room for improvement: how to keep the text view easier to clean without user intervention?
@@ 6.3. Physicality, Prototyping, Playfulness
This suggests that the design decision to give users interactive control via spatial direct manipulation in tandem with automation was successful; it enabled a form of playful exploration. This supports something akin to what professional designers and kids use to tinker with creative ideas. Note that P2 values ”help with making” rather than ”fully-automatic making,” meaning to them, interactivity was crucial for the playful prototyping feel.
@@ 6.4. Creative problem-solving
P1 was able to discover creative use of rule functionality by their own curiosity. In their case, they drew a boy, house, and tree. By creating two co-dependent rules, they wanted to see if they could recreate infinite loop functionality without using ”over and over” or ”forever.” Their solution was ”When boys collide with trees, boys move to houses” and ”When boys collide with houses, boys move to trees.”. This resulted in an infinite loop once the boy sketch collided with either the tree or the house. The example demonstrates the recombinability of the language-controllable primitives (Figure 16).
@@ 6.5. Re-applicability to other toolsets
In other words, users perceive DrawTalking as an interaction that is independent from a specific application. The suggestions for outputting code or plugging-into other applications with more advanced functionality hint at an interest in the way of controlling elements and a general excitement over how the ideas can be applied.
6.6. Naming by speech preferred to linking
We observed that labeling by speech (deixis) was unanimously preferred and used over linking and asked participants about their behavior. (See subsection 4.3 to recap labeling functionality.) It was considered more intuitive, direct, and similar to how people talk. On the other hand, pen+touch linking could be a useful fallback when freer grammar without directly-mentioned ”the” is preferred, e.g. when giving talks, or when referring to text far back in history without needing to speak again. Being able to link directly to a word could be more error-tolerant) P5 (an interactive computer graphics professor) gave such as example: ”I often talk about arrows in the context of vectors in my teaching”. Here, no deixis is used. Instead, the user would do a pen+touch operation between objects and the words ”arrows” and ”vectors” to do the labeling. This labeling can be performed in the background without requiring the speaker to use a specific grammar.
@@ 6.6. Naming by speech preferred to linking
6.7. Greater control for physical interfaces
At least one participant, P9 was a roboticist who could give specialized feedback. When asked about possibly incorporating into robotics P9 cautioned that for real-world environments, they might want to have even more control and self-specification of commands due to safety concerns, in spite of the loss of automation:
@@ 6.7. Greater control for physical interfaces
This suggests that our emphasis on user control was necessary and perhaps crucial for re-application of DrawTalking concepts into future use cases working with physical objects for the sake of safety. There is a clear trade-off between ”feeling magical” and ”feeling in-control.”
7. Discussion & Future Work
@@
Based on the primary contribution of our work (the interaction mechanism enabling sketching+sketching control) concrete implementations could use the concept to drive deeper levels of functionality and programmability than what is capable with our application prototype. The overarching idea, above all, is about exploring how to provide controllable feedback and facilitate a kind of vernacular style of programming, to extend our creative capabilities.
@@
There are many possible directions: Improving the system design and vocabulary; further-exploring visual-linguistic mappings; integrating with other applications; longer-term studies applying our approach to domain-specific creative or educational workflows; multi-user collaborations; application to spatial experiences for interaction with the real-world. We describe several future possibilities below:
@@
Integration with other applications. A DrawTalking-like sketching application could represent a prototyping/open-ended phase of the creative process. Future work can explore how to take the interaction with these rough named sketches and convert them for use in later-stage production elements, where goals have become more specific. There is also an opportunity to explore how to convert back and forth between different representations of objects and behaviors to support moving seamlessly between ideation and production phases. For example, we suspect that our strategy of naming objects independent of their representation could be useful as a way of generating scenes from the sketch representation.
@@
Semantics-visual mapping augmentation. There are several unexplored concepts relevant to semantics-visual mapping. For example, the use of sketches to represent abstract concepts could be evaluated as additional means for controlling content, in addition to the concrete sketches and language we’ve used. The semantics diagram, further, could be extended to support greater levels of editability. Another future area might involve context-sensitivity: What if the user’s culture and assumptions should produce different behaviors and visual-semantics mappings? How does DrawTalking, further, change when adapted to other languages?
Language input that is more natural, reducing user specification. This could potentially be addressed with a more robust and feature-complete back-end: e.g. improve the language processing implementation and/or integrate with faster generative language models for translation to simpler instructions.
Generative language models as of writing are too slow for interactive time, but in the future might be used to convert fully natural language into structured subsets easy for systems like ours to parse, or to generate commands based on application or scene context to reduce user-over-specification. This could allow for creation and management of larger persistent simulations.
@@
Spontaneous authoring of new primitives without programming. Exploring interactions for richer behavior-authoring interactions compatible our approach would be an excellent direction. We can imagine a number of possibilities: exploring extensions to the existing semantics diagram interface for lightweight editing; using DrawTalking’s controls to drive other applications with their own unique primitives; combining with the capabilities of generative models; crowd-sourcing programmed-functionality through shared libraries of domain-specific functionality; introducing multiple layers of programming capability in the same interface for different audiences (e.g. artists, programmers), trading-off simplicity for granular control as is common in many applications.
@@
Possibility for generalizability & scalability. We envision that combinations of DrawTalking-like functionality and different implementations and domain-specific applications could lead to a more-scalable ecosystem of interfaces. We also believe it may be useful to decouple the implementation of natural language processing from more deterministic processing, to facilitate reproducible research and support stabler interactive environments; in other words: stages of language compilation from probabilistic to deterministic instructions.
@@
Lastly, although we’ve focused on rough sketching, conceptually there is no limitation preventing sketches from being replaced with other forms of controllable digital or tangible objects. A sketches can simply reduce to a collections of variable properties and representations.
@@
We are excited to see ideas taken in new directions for playful and creative human-machine interfaces beyond our initial scope. We think that there’s an opportunity to seek-out more human-computer interactions that draw from our natural behavior for inspiration.

### v4

@@ 5.2. Processing Steps
The system receives language input and outputs a directed graph data structure encoding the dependencies between words and the words’ semantic roles.
Our implementation uses the Spacy (v 3.1.4)(Montani et al., 2023) library to output a dependency tree per-sentence, along with a library called ”coreferee” to fill-in coreference information. Accuracy of this phase is tied to the chosen NLP method, not our interface. Spacy tended to reproduce the same results for the same sentence structures, which made visual output predictable.
@@
Lastly, the application traverses the structure and generates a final application-defined structure S3 that it can evaluate – in our case, a mix between scriptable virtual machine and animation engine akin to (Perlin and Goldberg, 1996; Resnick et al., 2009) or a runnable node-based program (cycling74, 2018). S3 can call these scripts and retain S2 for reference. When traversing nodes in S2 labeled ACTION (the verbs) we lookup the appropriate previously-defined script for that action and insert a reference to it in S3. The arguments for that action, i.e. semantic role keys mapped to object ids or types, are inserted into a lookup table specific to the ACTION script instance so when the action executes, it knows what objects to modify. Loops and timed waits are also inserted into S3. The application evaluates the final structure according to its own interpretation. This results in any potential application-specific side-effect – in our case, animation, simulation, rule creation, application state-changes, and so on.
@@
Figure 12. Concrete Client/Server Details: Text data are sent to a server for natural language processing over a data channel and forwarded to an NLP-focused sub-process on the same machine. The semantic parse data return to the client. A command channel handles user events independently (e.g. discarding commands interrupts in-flight stale NLP).
@@
[A system diagram for the client-server architecture and data flow]A system diagram for the client-server architecture and data flow. The user is shown as a stick-figure with a quote bubble reading ”speech” and the text ”pen+touch” both connected to the main DrawTalking client node in the system diagram (indicating possible sources of user input.) The client node contains a real-life photo of the researcher’s iPad Pro used for the prototype, atop a blue square with the name ”DrawTalking” on it, as well as a small cartoon microphone icon. Below the square reads: ”Main C++, C; Platform: Obj C, Swift, Metal Graphics; On-Device Speech Recognition”. 3 arrows connect the client node to a server node. The arrows are referred to as ”Wireless Local Connection” (to the server. Arrow 1 (outgoing) is for text data. Arrow 2(incoming) is for the semantics parse. Arrow 3 (outgoing) is for control messages to the server; In the server node is a real-life photo of the Macbook Pro used for the server, above two nodes: left: the control server labeled NodeJS; right the Language Processing Sub-Process labeled Python 3. 2 arrows between server nodes (both surrounded by an enclosing rectangle): arrow 1: (outgoing) is for text. arrow 2: (incoming) is for the resulting semantics parse, which is sent back to the control server and back to the DrawTalking client along the other arrow incoming to the DrawTalking client.
@@
5.3. Current Limitations of Semantic Parsing
We focused on supporting narrative third-person form for this prototype due to our emphasis on scenarios involving explanation and storytelling. Passive tense, split infinitives, dangling prepositions, and idioms among others are unsupported. Other tenses and forms are supported as input, but behave the same as present tense. For example, past tense (”the dog jumped”), future tense (”the dog will jump”), present-progressive (”the dog is jumping”), and imperative (”dog, jump”) are equivalent to ”the dog jumps.” This allows for variety in the storytelling, but a future implementation might wish to distinguish between tenses for greater expressiveness.
Rules must sometimes be specified verbosely to avoid ambiguity in more complex cases. For example: ”When dogs collide with cats dogs jump and cats jump.” This could be expressed less verbosely, but the prototype at-present cannot disambiguate alternatives such as ”When dogs collide with cats they jump.” Here, ”they” is ambiguous because it could refer to ”dogs” only, ”cats” only, or as intended in this example, both. The first form is more verbose, but unambiguous. We support the second, but just as in English, the result might be unclear and the system will likely pick either ”dogs” or ”cats.”
Note that the semantics diagram still appears when a command is created from unsupported grammar. However, the diagram might show incorrect or missing object mappings and semantic roles. In such cases, the user can discard the malformed command or link missing objects to the diagram as in Figure 4. The user must learn and restrict themselves to the specific grammar to get the most consistent results. We cover how we might support more natural speech in section 8.
Although commands are user-initiated by design, a potential limitation is that commands must be created and confirmed one-by-one. Continuous speech input has unclear sentence boundaries, so we do not currently integrate a component to predict where a command should end because this could be unreliable and would take away control. To avoid rushing the user, we did not rely on a timeout either. We chose to give the user control over the sentence boundaries via the speech command button and interaction with the text transcript view (Figure 5). Different implementations could try automating commands without compromising user-control.
Additionally, the procedure to transform S1 into S2 (subsection 5.2) is coupled to the NLP library of choice by design. S2 and onwards are generic. An implementer would need to be aware of the need to build their own version of this translation step if their NLP back-ends use different dependency labels.
@@ 7. Results and Discussion
All participants understood and learned the mechanics, and produced thoughtful discussion. They tried their own ideas for how to make things or test their own understanding out of curiosity. All became comfortable with the core mechanics (controls, touch+deixis, commands, etc.) in 10-15 minutes. Some were excited to continue playing with the tool. They identified use cases including educational applications, rapid video game and design prototyping, paper prototyping, user interface design, presentation, language learning, and visual-oriented programming.
@@ 7.2. Semantics-visuals correspondence is valued
For the semantics diagram, participants understood it as a helpful means to introspect and correct the system’s understanding of user input. P2: ”[The diagram] is pretty important. Especially if you’re getting to used to the program, you might say something wrong and it interprets it in a different way and you can correct it.” P8: ”This is a really cool debug thing. I immediately understand what happens when I read it.” Additionally, the interface simplifies the input, making it easier to read: P7: ”[It’s] showing what it’s going to do. It’s taking the verbs and cutting all the other stuff out.” The feedback suggests that the diagram is valuable for instilling confidence in the system, and that the interface achieved D3.
For the transcript view, we observed users (e.g. P1, P5) habitually clearing the transcript using ”discard” on their own. We asked why and some participants reported wanting to keep the interface clean. We did not foresee this, but this user behavior serves a dual-purpose: the user is aware of what they say, which means they are perhaps likelier to catch errors, and they also help by keeping the system input from growing too long. We can take this as a positive: the system helps the user, and the user helps the system. We can also take this as room for improvement: how to keep the transcript view easier to clean without user intervention?
@@ 7.3. Physicality, Prototyping, Playfulness
This suggests that the design decision to give users interactive control via spatial direct manipulation in tandem with automation was successful; it enabled a form of playful exploration. This supports something akin to what professional designers and kids use to tinker with creative ideas. Note that P2 values ”help with making,” but doesn’t ask for full automation. To them, interactivity was important for the playful prototyping feel.
@@ 7.4. Creative problem-solving
P1 was able to discover creative use of rule functionality by their own curiosity. In their case, they drew a boy, house, and tree. By creating two co-dependent rules, they wanted to see if they could recreate infinite loop functionality without using ”over and over” or ”forever.” Their solution was ”When boys collide with trees, boys move to houses” and ”When boys collide with houses, boys move to trees.”. This resulted in an infinite loop once the boy sketch collided with either the tree or the house. The example demonstrates imaginative composition of language-controllable primitives (Figure 16).
@@ 7.5. Re-applicability to other toolsets
In other words, participants perceive DrawTalking as an interaction that is independent from a specific application. The suggestions for outputting code or plugging-into other applications with more advanced functionality hint at an interest in DrawTalking as a general control mechanism. Participants expressed excitement over how the ideas might be re-applied.
7.6. Observations and Feedback on Current Limitations of Language Understanding
Participants occasionally encountered some of the limitations of the semantic parsing described in subsection 5.3. For instance, P2 initially said, ”When the squirrel collides with the collider stop.” The prototype cannot currently infer that the rule should stop the squirrel. The subject (squirrel) needs to be re-specified as the target object to stop. The correct form would be: ”When the squirrel collides with the collider the squirrel stops.”. Similarly, P1 tried, ”The boy moves to the house and then to the tree.” The prototype does not infer the second subject and verb, so the input must be: ”The boy moves to the house and then the boy moves to the tree.” Other user errors included using imperative form without a subject (P5 ”Make a star”) (as imperatives currently require an explicit object), or expecting alternative behavior or synonyms (e.g. ”touch” as a substitute for ”collide with” or a variant of ”follow” that moves to a target only once, not continuously). Most of the unsupported cases suggest the following categories for improvement in future prototypes: reducing the need for user-specification, even in ambiguous or less-grammatically-correct cases; inferring implicit meaning from context; making the vocabulary and tenses more flexible. These are not tied to the DrawTalking concept, but could improve the experience of a concrete implementation.
As for the study experience, participants needed to become acquainted with the supported language structures. This was expected in the natural process of learning a new interface. In spite of the limitations, participants did not noticeably express frustration; they were excited. We suspect this is due to the following main factors:
Interaction design: invalid commands do not cause surprising or destructive side-effects (nothing happens) and redoing commands is fast. Pressing the discard button and speaking again does not incur too much of a time cost before the participant can see results. The world isn’t interrupted; it continues running and is interactive even while staging a command. In short, the quick and fluid error recovery might have helped reduce wait-time.
Participants were able to focus on the potential of the interactions and design: they were informed of the fact that they were being introduced to a prototype and they formulated responses with this in-mind. This is evidenced in the fact that each participant provided insight on potential improvements, integrations, and use cases assuming a more robust system. P1 and P5 especially, as seasoned computer scientists and educators, focused on the potential to integrate the workflow, feel, and functionality into other software.
This increases our confidence that participants focused their feedback on the concept rather than the implementation details.
7.7. Naming by speech preferred to text linking
Labeling by speech (deixis) was unanimously preferred to linking and was used most often or exclusively. (See subsection 4.3 to recap labeling functionality.) It was considered more intuitive, direct, and similar to how people talk. On the other hand, pen+touch linking could be useful when freer speech (without explicitly naming objects) is preferred, e.g. for flexibility while giving talks or to reference previous discussion without repeating words. Labeling by linking directly to a word instead of by speaking could allow for freer use of language, as the desired labels just need to be present in the narration. P5 (an interactive computer graphics professor) gave such as example: ”I often talk about arrows in the context of vectors in my teaching”. Here, no deixis is used. Instead, the user would perform a pen+touch operation between objects and the words ”arrows” and ”vectors” to do the labeling. The speaker can perform this labeling in the background without needing to use a specific grammar.
@@ 7.7. Naming by speech preferred to text linking
7.8. User control of physical elements and robots
Participant P9 was a roboticist who could give specialized feedback. When asked about possibly using DrawTalking for controlling robots, P9 said that for real-world environments, they might want to have even more control and self-specification of commands due to safety concerns, in spite of the loss of automation:
@@ 7.8. User control of physical elements and robots
This suggests that our emphasis on user control is likely important for reapplication of DrawTalking concepts in use cases involving physical objects. Additional safety considerations emerge and there are trade-offs between automation and control.
8. Future Work
@@
Based on the primary contribution of our work (the interaction mechanism enabling sketching+speaking control) concrete implementations could use the concept to drive deeper levels of functionality and programmability than what is capable with our application prototype. The overarching idea, above all, is about exploring how to provide controllable feedback and facilitate a kind of vernacular style of programming, to extend our creative capabilities.
@@
There are many directions: improving the system design and vocabulary; further-exploring visual-linguistic mappings; integrating with other applications; longer-term studies applying our approach to domain-specific creative or educational workflows; multi-user collaborations; application to spatial experiences for interaction with the real-world. We describe several future possibilities:
@@
Integration with other applications and extensibility. A DrawTalking-like sketching application could be a part of a prototyping/open-ended phase within a creative process. Future work can explore how to take the interaction with named rough sketches and convert them for use or replacement in later-stage production phases, in which goals have become more specific. The names and even the history of users’ direct manipulation and narration could be used to inform such a conversion process by providing context. We can also explore how to convert back and forth between different representations of objects and behaviors to support moving seamlessly between ideation and production phases. For example, we suspect that our strategy of naming objects independent of their representation could be useful as a way of generating scenes from the sketch representation.
Additionally, as P9 (section 6) suggests, DrawTalking could either provide animations to other tools, or else control external tools’ capabilities (not just limited animations) via plugin or remote API. This way, we treat DrawTalking as a controller rather than as a standalone application. (This also helps avoid vendor lock-in.) Vocabulary and functionality are potentially unlimited.
@@
Deeper Exploration of relationships and mappings between semantics and visuals. There are several unexplored concepts relevant to semantics-visual mapping. For example, the use of sketches to represent abstract concepts could be evaluated as additional means for controlling content, in addition to the concrete sketches and language we’ve used. The semantics diagram, further, could be extended to support greater levels of editability. Another future area might involve context-sensitivity: What if the user’s culture and assumptions should produce different behaviors and visual-semantics mappings? How might DrawTalking evolve when considering not only English, but also other languages, modes of communication, and cultural contexts?
Supporting language input that is more natural. Future work could work towards achieving increasingly natural and free-flowing speech input for interactive interfaces by implementing a more robust and feature-complete language processing and multi-modal input back-ends. We suspect that a translation layer from natural, more fragmented speech could convert input into simpler, structured, and deterministic instructions similar to what our interface uses. This direction could allow existing interfaces with deterministic instructions like ours to accept more expressive, less restricted language input, independent of their functionality. This could lead to reduced user-specification and a better approximation fully ”natural” language interfaces.
Making trade-offs: reasons why future interface designers might not want to use or rely on generative-artificial-intelligence models (e.g. LLMs), and why we did not use them. To achieve fully natural input, a suggestion might be to use generative AI models such as large language models (LLMs) to perform the aforementioned proposed translation of language input. These models could also generate visual content alongside user-drawn sketches. (A number of study participants made this suggestion.)
However, as per our design goals for this interface (subsection 3.3), the system must be interactive and instant, and the user must be in-control of the content’s representation. We considered LLMs, but they contradicted our design goals: reducing fluidity, user control over content, and system transparency and understandability. Interface designs with similar criteria might also encounter the same challenges and opt to make trade-offs.
LLMs are currently too slow to provide instant feedback, even if we use LLMs only to transform natural input into structured commands as we suggest. (This also raises additional research questions: How well can LLMs rewrite users’ speech and reflect intent?) In contrast, the NLP technology we chose responds nearly instantly.
Also, LLMs often exhibit unpredictable or unreproducible behavior, whereas the chosen NLP returns the same results for the same language input. We wanted determinism so command outcomes would be easy to reason about, especially to keep the scope of possibility manageable for the user study. Not using LLMs allowed for feasibility and stable development.
Still, generative-AI is promising given the right context and requirements. Future work could see how LLMs might be complementary to other methods (traditional NLP, procedural generation, and others). We believe a good direction for research would be to find the balance between the many approaches and solutions we have developed over time.
@@
Spontaneous authoring of new primitives without programming. Exploring interactions for richer behavior-authoring interactions compatible with our approach would be another direction. We can imagine a number of possibilities: exploring extensions to the existing semantics diagram interface for lightweight editing; using DrawTalking’s controls to drive other applications with their own unique primitives; combining procedural methods with the capabilities of generative models by finding the right balance; crowd-sourcing programmed-functionality through shared libraries of domain-specific functionality; introducing multiple layers of programming capability in the same interface for different audiences (e.g. artists, programmers), trading-off simplicity for granular control as is common in many applications.
@@
Possibilities for generalizability and scalability. We envision that combinations of DrawTalking-like functionality and different implementations and domain-specific applications could lead to a more-scalable ecosystem of interfaces. We also believe that by decoupling natural language processing implementations from the interface (rather than building directly around specific NLP implementations or probabilistic results) it might be easier to support reproducible research within interactive interface design. This is in-part why we chose to compile NLP information into intermediary, generic command structures. These could feasibly be used in future research as a compilation target by any other NLP or alternative approach without relying on a specific implementation. For example, this is where a gen-AI model could be explored without changing the rest of the system.
@@
Lastly, although we’ve focused on rough sketching, conceptually there is no limitation preventing sketches from being replaced with other forms of controllable digital or tangible objects. Sketches can represent collections of variables and properties mapped to some objects, whatever forms they might take.
@@
We are excited to see ideas taken in new directions for playful and creative human-machine interfaces beyond our initial scope. We think that there is an opportunity to seek-out new human-computer interactions that draw from our natural behavior for inspiration.